# Trapping an Atomic Ion without Dedicated Digital-to-analog Converters

### v1

@@ Abstract
Independent control of numerous electrodes in quantum charge-coupled device architectures presents a significant challenge for wiring and hardware scalability. To address this issue, we demonstrate a voltage control method based on time-division multiplexing (TDM). This approach utilizes a single high-update-rate digital-to-analog converter (DAC) to sequentially generate control signals for multiple electrodes, thereby reducing both the number of required DACs and associated wiring. We experimentally validate this concept by developing a 10-channel system that operates with only two DACs. The developed TDM-based voltage control system is applied to a surface-electrode trap, where we successfully trap a single ion. This approach offers a resource-efficient and scalable solution for advanced quantum computing systems based on trapped ions.
@@
In this work, we present an experimental proof of concept for a TDM-based approach to scalable electrode control. We develop a system capable of controlling ten independent trap electrodes using only two DACs, and validate its functionality by trapping a single ion in a surface-electrode trap. This result demonstrates the potential of TDM-based scalable electrode control as a practical solution for reducing classical hardware overhead in large-scale QCCD systems.
@@
A likely cause of this deviation is an offset in the input voltage . The measured [dashed curve in Fig.3(b)] spans from 0.26 V to 2.79 V, deviating from the ideal range of 0 V to 2.5 V. Substituting these measured values into Eq. (1), the corresponding is calculated to range from -7.9(3) V to 15.4(5) V. The uncertainties represent values, derived by propagating the tolerances of and under a uniform distribution 222 and are general-purpose chip resistors, RK73B1JTTD822J and RK73B1JTTD102J, respectively, manufactured by KOA Speer Electronics, Inc. According to their datasheet, both resistors have a tolerance of . . The results are consistent with the presence of an input offset, suggesting it is the primary contributing factor. Although the origin of the offset is presently unknown, it is expected to be correctable through circuit modification.
@@
In summary, we have demonstrated the successful trapping of a single ion using the proposed TDM-based voltage control scheme. This method enables the control of ten DC electrodes in a surface trap using only two DACs, thereby eliminating the need for ten independent voltage sources. While the current study establishes a functional proof of concept, several challenges must be addressed for future implementation and system scaling.

### v2

# Trapping an Atomic Ion using Time-Division Multiplexed Digital-to-Analog Converters

@@ Abstract
Independent control of numerous electrodes in quantum charge-coupled device architectures presents a significant challenge for wiring and hardware scalability. To address this issue, we demonstrate a voltage control method based on time-division multiplexing (TDM). This approach utilizes a single high-update-rate digital-to-analog converter (DAC) to sequentially generate control signals for multiple electrodes, thereby reducing both the number of required DACs and associated wiring. We experimentally validate this concept by developing a 10-channel system that operates with only two DACs. The developed TDM-based voltage control system is applied to a surface-electrode trap, where we successfully trap a single ion and demonstrate a simple ion transport primitive. This approach offers a resource-efficient and scalable solution for advanced quantum computing systems based on trapped ions.
@@
In our previous work [12], we proposed a TDM-based control scheme and demonstrated only a preliminary proof-of-concept with five multiplexed channels, without application to an actual ion trap. In this work, we present an experimental proof of concept for a TDM-based approach to scalable electrode control. We develop a system capable of controlling ten independent trap electrodes using only two DACs, and validate its functionality by trapping a single ion in a surface-electrode trap. We also demonstrate a simple ion transport primitive. This result demonstrates the potential of TDM-based scalable electrode control as a practical solution for reducing classical hardware overhead in large-scale QCCD systems.
@@
A likely cause of this deviation is an offset in the input voltage . The measured [dashed curve in Fig.3(b)] spans from 0.26 V to 2.79 V, deviating from the ideal range of 0 V to 2.5 V. Substituting these measured values into Eq. (1), the corresponding is calculated to range from -7.9(3) V to 15.4(5) V. The uncertainties represent values, derived by propagating the tolerances of and under a uniform distribution 222 and are general-purpose chip resistors, RK73B1JTTD822J and RK73B1JTTD102J, respectively, manufactured by KOA Speer Electronics, Inc. According to their datasheet, both resistors have a tolerance of . . The results are consistent with the presence of an input offset, suggesting it is the primary contributing factor. We find that the offset originates from the regulator that supplies the reference voltage for the first-stage OP-amp (AD8099), which is designed to produce 0–2.5 V under nominal conditions. The non-inverting input of this OP-amp is tied to a reference voltage obtained by resistive division of the regulator output, nominally 0.80 V. However, the regulator does not operate properly and outputs 5.43 V; the reference voltage rises to 0.89 V. Consequently, the output range shifts from the intended 0–2.5 V to approximately 0.28–2.78 V, consistent with the measured value of 0.26–2.79 V.
@@
Furthermore, shuttling of a single trapped ion is performed by applying a time-dependent voltage generated by the TDM-based control system. The experimental conditions differ from those used for static trapping of a single ion. In this experiment, the RF confinement voltage is set to 240 V, and the vacuum pressure is 4.1 Pa. A qCMOS camera (C15550-20UP, Hamamatsu Photonics) is used to record the ion images. The result of the shuttling experiment is shown in Fig. 4(c). The image shows the ion being transported along the -axis direction when the time-dependent voltage is applied to the side DC electrodes. The applied waveform consists of eleven discrete voltage steps, each updated every 500 ms. The waveform for the ion transport is obtained using the numerical method described in Ref. [21]. Note that the voltage update rate is relatively slow; the shuttling speed is not optimized in this demonstration. Here, we simply show that the shuttling primitive can be realized using the TDM-based voltage control scheme. The center DC electrode is held at a fixed bias of 0.24 V. During the shuttling operation, both the 397 nm and 866 nm laser beams are kept continuously on along the -axis.
In summary, we have demonstrated the successful trapping and shuttling of a single ion using the proposed TDM-based voltage control scheme. This method enables the control of ten DC electrodes in a surface trap using only two DACs, thereby eliminating the need for ten independent voltage sources. While the current study establishes a functional proof of concept, several challenges must be addressed for future implementation and system scaling.
@@
Here, we discuss the feasibility of introducing the TDM-based voltage control scheme into a cryogenic environment, taking the 40 K stage as an example. The dominant source of power dissipation is the final-stage OP-amp. For the THS4631 used in this work, the quiescent current is 12.5 mA, giving a static dissipation of 375 mW at V voltage supply. With a cooling power of W at the 40 K stage of a commercial Gifford–McMahon cryocooler, this would allow integration of about 100 channels. This estimate assumes a high-slew-rate amplifier (1000 V/s). However, such specifications are not strictly required for ion-trap operation, and substantially lower-slew-rate devices may be considered. For example, the LT1351 (Analog Devices, Inc., 200 V/s slew rate) exhibits a quiescent current of only 0.25 mA, enabling integration of 5000 channels. Further scalability may be achieved with the architecture shown in Fig. 5, where signals are pre-amplified at a higher-temperature stage. In this configuration, switches must tolerate voltage amplitudes of approximately 10 V and remain operable at cryogenic temperatures. Although the attainable switching speed would be limited, the need to place OP-amps directly inside the cryostat is eliminated, thereby reducing the thermal load. If suitable cryogenic switches operable at 4 K become available, realization of this scheme at the 4 K stage would also be plausible.